5. Prove that rb b² – a² xdx = = 2 (Hint: Evaluate the corresponding limit of Riemann sum.)”
Answer #1

à ad n a da let n be the number of intervals then interval width (ox= 6-9). with Dr = ba from plato (Hab , we have followingI) 15 from him & [469+09:3 2 = dice la com fotby20] we have E > + A > 1 limon [ a 6-a). (n) Concour] In this com [a66-) 4 (-)